Yesterday Sandi sold 1,000 shares of stock that she owned for $55 per share. When she purchased the stock two years ago, Sandi paid $60 per share. Every three months during the time that she held the stock, Sandi received a quarterly dividend equal to $0.75 per share. A total of eight dividends were received. (a) What return (yield) did Sandi earn during the two years she held the stock? (b) If the price of the stock was $55 per share one year ago, what return did Sandi earn in each year she held the stock?
Answer #1

Solution-

a) Total dividend = 0.75*8= $6

Holding period return = ( selling price - buying price + dividend)/buying price

Holding period return = (55-60+6)/60= 0.0167 or 1.67%

b) Return in year 1 :-

Holding period return = (55-60+0.75*4)/60= 3.33%

Return in year 2:-

Holding period return = (55-55+0.75*4)/55= 0.0545 or 5.45%